OpenStudy (anonymous):

Find an equation of a line parallel to y=5x+5 and passing through the point (2, 5). y=

OpenStudy (anonymous):

y=5x-5

OpenStudy (anonymous):

Thank you so much. I really appreciate it. Can you show me how you solved that for future reference? I have a quiz coming up soon, and I would love to learn it.

OpenStudy (anonymous):

find the slope of the equation that is y=5x+5 let it be y=mx+b (m is the slop). then apply point slope form into the point (2,5). (y-5)=m(x-2)

OpenStudy (anonymous):

Oh okay thank you :)